A stock's returns have the following distribution:

Demand for the
Company's Products
Probability of This
Demand Occurring
Rate of Return If
This Demand Occurs
Weak 0.1 (46%)
Below average 0.3 (10)   
Average 0.3 11  
Above average 0.1 29  
Strong 0.2 66  
1.0

Assume the risk-free rate is 4%. Calculate the stock's expected return, standard deviation, coefficient of variation, and Sharpe ratio. Do not round intermediate calculations. Round your answers to two decimal places.

Stock's expected return:   %

Standard deviation:   %

Coefficient of variation:

Sharpe ratio:
Answer #1

Expected Return = [0.1(-0.46) + 0.3(-0.10) + 0.3(0.11) + 0.10(0.29) + 0.2(0.66)]

Expected Return = 0.118

Standard Deviation = [(0.1(-0.46 - 0.118)2 + 0.3(-0.10 - 0.118)2 + 0.3(0.11 - 0.118)2 + 0.1(0.29 - 0.118)2 + 0.2(0.66 - 0.118)2]1/2

Standard Deviation = 0.3308

CV = 0.3308/0.118 = 280.34%

Sharpe Ratio = (0.118 - 0.04)/0.3308

Sharpe Ratio = 0.24